64.0 mL of a 1.70 M solution is diluted to a total volume of 248 mL. A 124-mL portion of that solution is diluted by adding 195 mL of water. What is the final concentration? Assume the volumes are additive.
Answer #1

By using dilution formula.

M1V1 = M2V2

M1 = molarity of concentrated solution = 1.7 M

V1 = Volume of concentrated solution. = 64 ml

M2 = molarity of dilute solution. = ?

V2 = Volume of dilute solution. = 248ml

M2 = M1V1/V2 = 64*1.7/248 = 0.4387 M

now 124-mL portion of that solution is diluted by adding 195 mL of water

M1 = molarity of concentrated solution = 0.4387 M

V1 = Volume of concentrated solution. = 124 ml

M2 = molarity of dilute solution. = ?

V2 = Volume of dilute solution. = 124+195 = 319 ml

M2 = M1V1/V2 = 124*0.4387/319 = 0.1705 M